# Exponentially Weighted Instance-Aware Repeat Factor Sampling for Long-Tailed Object Detection Model Training in Unmanned Aerial Vehicles Surveillance Scenarios

### v1

@@ Abstract
Object detection models often struggle with class imbalance, where rare categories appear significantly less frequently than common ones. Existing sampling-based rebalancing strategies, such as Repeat Factor Sampling (RFS) and Instance-Aware Repeat Factor Sampling (IRFS), mitigate this issue by adjusting sample frequencies based on image and instance counts. However, these methods are based on linear adjustments, which limit their effectiveness in long-tailed distributions. This work introduces Exponentially Weighted Instance-Aware Repeat Factor Sampling (E-IRFS), an extension of IRFS that applies exponential scaling to better differentiate between rare and frequent classes. E-IRFS adjusts sampling probabilities using an exponential function applied to the geometric mean of image and instance frequencies, ensuring a more adaptive rebalancing strategy. We evaluate E-IRFS on a dataset derived from the Fireman-UAV-RGBT Dataset and four additional public datasets, using YOLOv11 object detection models to identify fire, smoke, people and lakes in emergency scenarios. The results show that E-IRFS improves detection performance by 22% over the baseline and outperforms RFS and IRFS, particularly for rare categories. The analysis also highlights that E-IRFS has a stronger effect on lightweight models with limited capacity, as these models rely more on data sampling strategies to address class imbalance. The findings demonstrate that E-IRFS improves rare object detection in resource-constrained environments, making it a suitable solution for real-time applications such as UAV-based emergency monitoring.
@@ I Introduction
These contributions refine class imbalance handling in object detection. E-IRFS enhances rare-class representation without excessive oversampling, benefiting lightweight models that struggle with imbalance. This makes it well-suited for UAV-based monitoring and edge applications. Table I summarizes the symbols and notations used throughout the paper for consistency.
@@
$r_{i}$ Repeat factor for image $i$
@@
$p_{i}$ Probability of selecting image $i$ for training
@@ II-B2 Instance-Aware Repeat Factor Sampling (IRFS)
$r_{c}=\max\left(1,\sqrt{\frac{t}{\sqrt{f_{i,c}\cdot f_{b,c}}}}\right)$ (2)
@@ III-A Exponentially Weighted Instance-Aware Repeat Factor Sampling (E-IRFS)
$r_{c}=\exp\left(\alpha\cdot\sqrt{\frac{t}{\sqrt{f_{i,c}\cdot f_{b,c}}}}\right)$ (3)
@@ III-B Image-Level Sampling Using E-IRFS
$p_{i}=\frac{r_{i}}{\sum_{j}r_{j}}\vspace{-2mm}$ (4)
where, $\sum_{j}r_{j}$ computes the total sum of all image-level repeat factors across the entire dataset and $p_{i}$ represents the normalized probability of selecting image ii for training. This ensures that the sampling probabilities sum to 1, maintaining training stability while emphasizing rare categories effectively.
@@ III-C4 Effect on Convergence and Sampling Stability
Since the exponential function is always positive and the second term is negative,$r_{c}$ is monotonically decreasing. This ensures that as the image frequency $f_{i,c}$ decreases, the sampling rate increases, reinforcing rare-class representation in training. Furthermore, the second derivative is as follows:
@@ III-D Benchmark Database
We created a custom benchmark for object detection by merging five distinct datasets ([1], [11], [12], [13], [14]) and intentionally selecting a subset of the total images from the merged datasets to evaluate our methodologies. This subset was carefully chosen to ensure a diverse yet imbalanced distribution for evaluation purposes. This intentional class imbalance in the subset provides a challenging scenario for evaluating fire, smoke, human presence, and water body detection, fostering robust model training under diverse conditions. In our custom dataset, we have 40,384 training images and 11,953 validation images. The validation set is constructed by concatenating the official validation splits from the five merged datasets rather than a manually selected subset.
@@ III-E Evaluation Metrics and Protocols
The effectiveness of the YOLOv11 object detection models was evaluated using standard metrics provided by the Ultralytics validation system. These metrics assess the accuracy and reliability of the model in identifying objects across different conditions. To quantify detection performance, we used mean Average Precision at an Intersection over Union (IoU) threshold of 0.5 (mAP50) and mean Average Precision across multiple IoU thresholds ranging from 0.5 to 0.95 in steps of 0.05 (mAP50-95). The mAP50 metric evaluates detection performance by computing the precision-recall curve at a fixed IoU threshold of 0.5, allowing for some localization tolerance. In contrast, mAP50-95 provides a more strict assessment by averaging precision across multiple IoU thresholds, thereby offering a more comprehensive measure of detection accuracy and localization robustness. These metrics facilitate a standardized comparison of model performance across different object scales and levels of complexity within the dataset, ensuring a rigorous evaluation of the proposed approach.
@@ IV-B Performance Comparison of Re-Balancing Methods
To further assess the effect of re-balancing with data augmentation, Table III presents results with augmentation applied across all methods, including the baseline. While RFS and IRFS improve mAP5050{50}50 by 11%, their gains remain marginal compared to training without augmentation. In contrast, E-IRFS achieves a 22% increase in mAP5050{50}50 and a 13% improvement in mAP50-95, further demonstrating its advantage in addressing class imbalance, as shown in Figure 2. These results suggest that while augmentation benefits all methods, E-IRFS benefits the most due to its exponential weighting, which amplifies the effect of balanced sample exposure. The performance differences highlight a key limitation of RFS and IRFS, where their linear weighting does not sufficiently enhance rare-class representation.
@@ IV-C Threshold and Scaling Factor Analysis in E-IRFS
For example, for the small model YOLOv11-Nano (Table IV), setting $\alpha=2.0$ with $t=0.0001$ yields a 22% improvement in mAP50, outperforming lower $\alpha$ values where the effect of exponential scaling is weaker. A similar trend is observed in YOLOv11-Large, where the same configuration leads to an increase 18%. Lower values of $\alpha$, such as 0.5 or 1.0, result in more conservative adjustments, which reduce the ability to compensate for extreme class imbalances. The improvement at $\alpha=2.0$ and a low threshold is due to the exponential function in Equation 4. When the geometric mean is small, indicating a rare class, exponentiation amplifies the repeat factor more significantly, ensuring stronger sampling weights and better representation in training.
@@ IV-D Impact of Re-Balancing Methods on Class-Wise Performance
For the Lake category, mAP5050{50}50 increased from 0.02 to 0.09, corresponding to a relative improvement of 350%, while mAP50955095{50-95}50 - 95 increased from 0.009 to 0.032, reflecting a gain of 255%. Similarly, the Fire class showed an increase of 65% in mAP5050{50}50 and 31% in mAP50955095{50-95}50 - 95. The Human class benefited from an increase 16% in mAP5050{50}50 and an 18% increase in mAP50955095{50-95}50 - 95, indicating balanced improvements. In contrast, the Smoke class exhibited only marginal improvements, suggesting that it was already well-represented in the dataset.
@@ V Conclusion
This paper introduces E-IRFS, a novel rebalancing strategy designed to improve rare-class detection in long-tailed object detection. The experimental results demonstrate that E-IRFS outperforms existing methods, n detecting rare objects such as fire and water bodies in UAV-based monitoring. E-IRFS is especially beneficial for lightweight models with fewer parameters, such as YOLOv11-Nano, since they lack the capacity to learn long-tailed distributions effectively, making them more dependent on data sampling strategies. E-IRFS benefits from a lower threshold and a higher scaling factor, allowing more adaptive sampling adjustments. Beyond its empirical improvements, it provides a structured approach to addressing class imbalance, with differentiation between frequent and rare categories, while highlighting the importance of adaptive sampling strategies. Future work may explore extending E-IRFS to other recognition tasks and integrating it with additional augmentation techniques to enhance its performance in highly imbalanced datasets.

### v2

@@ Abstract
Object detection models often struggle with class imbalance, where rare categories appear significantly less frequently than common ones. Existing sampling-based rebalancing strategies, such as Repeat Factor Sampling (RFS) and Instance-Aware Repeat Factor Sampling (IRFS), mitigate this issue by adjusting sample frequencies based on image and instance counts. However, these methods are based on linear adjustments, which limit their effectiveness in long-tailed distributions. This work introduces Exponentially Weighted Instance-Aware Repeat Factor Sampling (E-IRFS), an extension of IRFS that applies exponential scaling to better differentiate between rare and frequent classes. E-IRFS adjusts sampling probabilities using an exponential function applied to the geometric mean of image and instance frequencies, ensuring a more adaptive rebalancing strategy. We evaluate E-IRFS on a dataset derived from the Fireman-UAV-RGBT Dataset and four additional public datasets, using YOLOv11 object detection models to identify fire, smoke, people and lakes in emergency scenarios. The results show that E-IRFS improves detection performance by 22% over the baseline and outperforms RFS and IRFS, particularly for rare categories. The analysis also highlights that E-IRFS has a stronger effect on lightweight models with limited capacity, as these models rely more on data sampling strategies to address class imbalance. The findings demonstrate that E-IRFS improves rare object detection in resource-constrained environments, making it a suitable solution for real-time applications such as UAV-based emergency monitoring. The code is available at: https://github.com/futurians/E-IRFS.
@@ I Introduction
These contributions refine class imbalance handling in object detection. E-IRFS enhances rare-class representation without excessive oversampling, benefiting lightweight models that struggle with imbalance. This makes it well-suited for UAV-based monitoring and edge applications. Table I summarizes the symbols and notations used throughout the paper for consistency. The code for E-IRFS and the implementation details of the custom benchmark dataset are available at: https://github.com/futurians/E-IRFS.
@@
$r_{i}$ Repeat factor for image ii
@@
$p_{i}$ Probability of selecting image ii for training
@@ II-B2 Instance-Aware Repeat Factor Sampling (IRFS)
$r_{c}=\max\left(1,\sqrt{\frac{t}{\sqrt{f_{i,c}\cdot f_{b,c}}}}\right),$ (2)
@@ III-A Exponentially Weighted Instance-Aware Repeat Factor Sampling (E-IRFS)
$r_{c}=\exp\left(\alpha\cdot\sqrt{\frac{t}{\sqrt{f_{i,c}\cdot f_{b,c}}}}\right),$ (3)
@@ III-B Image-Level Sampling Using E-IRFS
$p_{i}=\frac{r_{i}}{\sum_{j}r_{j}},\vskip-5.69054pt$ (4)
where $\sum_{j}r_{j}$ computes the total sum of all image-level repeat factors across the entire dataset and $p_{i}$ represents the normalized probability of selecting image ii for training. This ensures that the sampling probabilities sum to 1, maintaining training stability while emphasizing rare categories effectively.
@@ III-C4 Effect on Convergence and Sampling Stability
[…]
@@ III-D Benchmark Database
We created a custom benchmark for object detection by merging five different datasets ([1], [11], [12], [13], [14]). To construct the training and validation sets of our custom dataset, we used the original training and validation splits from four of these datasets [11, 12, 13, 14] without applying any selection or filtering. From the dataset in [1], we selectively used only the Unimodal-RGB subset from the Multimodal collection, excluding the thermal modality. Furthermore, we included only a curated set of video frames that contained at least one of the following objects: fire, smoke, humans, or lakes. The resulting custom dataset comprises 40,384 training images and 11,953 validation images. The validation set was formed by concatenating the official validation splits from the five merged datasets rather than selecting a separate subset. This setup ensures a diverse, yet intentionally unbalanced distribution for evaluation purposes. The class imbalance provides a challenging scenario for evaluating fire, smoke, human presence, and water body detection, fostering robust model training under diverse conditions. We also evaluate E-IRFS on two well-known balanced datasets, CIFAR-10 [15] and Caltech-101 [16], to assess its behavior when class distributions are uniform. This allows us to verify whether the exponential reweighting introduces performance degradation outside long-tailed scenarios.
@@ III-E Evaluation Metrics and Protocols
The effectiveness of the YOLOv11 object detection models was evaluated using standard metrics provided by the Ultralytics validation system. These metrics assess the accuracy and reliability of the model in identifying objects across different conditions. To quantify detection performance, we used mean Average Precision at an Intersection over Union (IoU) threshold of 0.5 (mAP50) and mean Average Precision across multiple IoU thresholds ranging from 0.5 to 0.95 in steps of 0.05 (mAP50-95). The mAP50 metric evaluates detection performance by computing the precision-recall curve at a fixed IoU threshold of 0.5, allowing for some localization tolerance. In contrast, mAP50-95 provides a more strict assessment by averaging precision across multiple IoU thresholds, thereby offering a more comprehensive measure of detection accuracy and localization robustness. These metrics facilitate a standardized comparison of model performance across different object scales and levels of complexity within the dataset, ensuring a rigorous evaluation of the proposed approach. For classification tasks on CIFAR-10 and Caltech-101, we report Top-1 and Top-5 accuracy, which measure whether the correct class is the model first or among its five highest-ranked predictions. These metrics are standard for evaluating classification without localization.
@@ IV-B Performance Comparison of Re-Balancing Methods
To further assess the effect of re-balancing with data augmentation, Table III presents results with augmentation applied across all methods, including the baseline. While RFS and IRFS improve mAP50{50} by 11%, their gains remain marginal compared to training without augmentation. In contrast, E-IRFS achieves a 22% increase in mAP50{50} and a 13% improvement in mAP50-95, further demonstrating its advantage in addressing class imbalance, as shown in Figure 2. These results suggest that while augmentation benefits all methods, E-IRFS benefits the most due to its exponential weighting, which amplifies the effect of balanced sample exposure. The performance differences highlight a key limitation of RFS and IRFS, where their linear weighting does not sufficiently enhance rare-class representation.
@@ IV-C Threshold and Scaling Factor Analysis in E-IRFS
For example, for the small model YOLOv11-Nano (Table IV), setting $\alpha=2.0$ with $t=0.0001$ produces a 22% improvement in mAP50, outperforming lower $\alpha$ values where the effect of exponential scaling is weaker. A similar trend is observed in YOLOv11-Large, where the same configuration leads to an increase 18%. Lower values of $\alpha$, such as 0.5 or 1.0, result in more conservative adjustments, which reduce the ability to compensate for extreme class imbalances. The improvement at $\alpha=2.0$ and a low threshold is due to the exponential function in Equation 4. When the geometric mean is small, indicating a rare class, exponentiation amplifies the repeat factor more significantly, ensuring stronger sampling weights and better representation in training.
@@ IV-D Impact of Re-Balancing Methods on Class-Wise Performance
For the Lake category, mAP50{50} increased from 0.02 to 0.09, corresponding to a relative improvement of 350%, while mAP5095{50-95} increased from 0.009 to 0.032, reflecting a gain of 255%. Similarly, the Fire class showed an increase of 65% in mAP50{50} and 31% in mAP5095{50-95}. The Human class benefited from an increase 16% in mAP50{50} and an 18% increase in mAP5095{50-95}, indicating balanced improvements. In contrast, the Smoke class exhibited only marginal improvements, suggesting that it was already well-represented in the dataset.
@@
IV-E Generalization to Balanced Datasets
To evaluate the applicability of E-IRFS beyond long-tailed scenarios, we tested it on two balanced datasets: CIFAR-10 [15] and Caltech-101 [16]. These experiments aim to determine whether the exponential weighting mechanism introduces performance degradation or bias when class distributions are approximately uniform. We trained the YOLOv11-Nano classification model on both datasets using baseline and E-IRFS configurations with $\alpha=0.5$, $\alpha=1.0$, and $\alpha=2.0$, fixing the threshold at $t=0.1$.
TABLE IX: Accuracy on Balanced Datasets (Top-1 / Top-5)
Dataset Method Top-1 Acc Top-5 Acc
CIFAR-10 Baseline 0.802 0.989
E-IRFS ($\alpha=0.5$, $t=0.1$) 0.802 0.989
E-IRFS ($\alpha=1.0$, $t=0.1$) 0.802 0.989
E-IRFS ($\alpha=2.0$, $t=0.1$) 0.802 0.989
Caltech-101 Baseline 0.920 0.986
E-IRFS ($\alpha=0.5$, $t=0.1$) 0.927 0.986
E-IRFS ($\alpha=1.0$, $t=0.1$) 0.924 0.985
E-IRFS ($\alpha=2.0$, $t=0.1$) 0.920 0.984
As shown in Table IX, E-IRFS maintains performance parity with the baseline across all settings on CIFAR-10 and Caltech-101. Minor fluctuations in Caltech-101 are within acceptable bounds and do not suggest systematic degradation. These results confirm that E-IRFS does not introduce bias or degrade performance on balanced datasets, supporting its robustness and generalization beyond imbalanced scenarios.
@@ V Conclusion
This paper introduces E-IRFS, a novel rebalancing strategy designed to improve rare-class detection in long-tailed object detection. The experimental results demonstrate that E-IRFS outperforms existing methods in detecting rare objects such as fires and water bodies in UAV-based monitoring. E-IRFS is especially beneficial for lightweight models with fewer parameters, such as YOLOv11-Nano, since they lack the capacity to learn long-tailed distributions effectively, making them more dependent on data sampling strategies. E-IRFS benefits from a lower threshold and a higher scaling factor, allowing more adaptive sampling adjustments. Beyond its empirical improvements, it provides a structured approach to addressing class imbalance, with differentiation between frequent and rare categories, while highlighting the importance of adaptive sampling strategies. E-IRFS shows sensitivity to selection of $\alpha$ and $t$, but our results indicate that effective values can generalize across datasets with minimal tuning. Future work may explore the extension of E-IRFS to other recognition tasks and the integration of it with additional augmentation techniques to improve its performance in highly imbalanced datasets. Furthermore, a critical direction involves investigating methodologies to determine the optimal $\alpha$ and $t$ parameters based on statistical priors from new datasets. This would reduce the need for extensive manual tuning, thus enhancing the practicality and applicability of the method across diverse domains.
Acknowledgment
This research was supported by the Research Council of Finland 6G Flagship program (Grant Number 369116), FIREMAN project (Grant Number 348008), and Business Finland Neural Publish Subscribe for 6G Project (diary number 8754/31/2022).